dr - 0 dz 1 dx P x² + y² + 하공 8. OP 11 위공 d츠 1 ㅇ d² 0= 이1) Here are two rank 1 tensors of differing types, written in Cartesian co-ordinates: 1 -4 vi(x, y, z) = wi(x,y,z) 2 3 a) Cal
Answer #1

(a) The scalar product of the given two tensors can be obtained as

2 2 هوا + 2 مهر ماه اوه يه = نوه خود ہے (- ) - ( + 4(- 4) + 3.1 _ 6 - = 3+ و- = 3 + 8 - 1 - =.

(b) The Jacobian for transformation of Cartesian coordinates into cylindrical polar coordinates is given by

т Отә O2 Әт Эк — Әу Ә. ох де Әу 22 Ә2 Ә2 Ә 22 ә2 Әу ( ) т ут х» үр = Jху O — у/ О О o x=2 Cose Cos o — Sino Cosa 0 9 = msine.

Hence the transformed tensors can be obtained as

( رو رع) ذیه ک = (2 ر2 ,2) اما Cos o Sino - 4 :) -Sino Cose O - Coso - 4 Sind Sing - 4 Cos

and

.

(c) The dot product in the cylindrical polar coordinates is

re; o = (- Cose +2 Sino) (Cos 0 - 9Sino) +(+ Sino +2 Cos 0) (-sino-4 los e) +3 Cos² + 2 sine los at 4 sino lose = -8 Sin²o -

which is same as the previous one.

(d) In cylindrical polar coordinates,

По а пор,

where g^{\alpha \beta } is the metric tensor in cylindrical polar coordinates.